# Eclipsed X-ray Bursts from Magnetar SGR J1935+2154 and the Fireball Measurements

### v2

@@ II.2 Eclipse Modeling
$\chi$ (deg) [0180][0-180] $17.46_{-6.87}^{+8.99}$ $17.11_{-8.94}^{+9.69}$ $18.54_{-8.16}^{+15.65}$ $8.04_{-3.56}^{+5.39}$
$\xi$ (deg) [0180][0-180] $155.92_{-11.71}^{+9.35}$ $162.22_{-9.20}^{+7.31}$ $149.38_{-23.41}^{+13.22}$ $172.00_{-5.43}^{+3.50}$
$\phi$ (deg) [0360][0-360] $343.75_{-1.00}^{+1.09}$ $297.60_{-0.32}^{+0.33}$ $316.85_{-0.69}^{+0.66}$ $274.27_{-1.21}^{+1.15}$
$d/R_{\mathrm{ns}}$ [150][1-50] $5.38_{-1.58}^{+3.19}$ $7.65_{-2.53}^{+5.01}$ $8.23_{-3.15}^{+5.89}$ $7.96_{-3.09}^{+6.07}$
$F_{\mathrm{p}}$ (ph cm-2 s-1) [11000][1-1000] $727.62_{-1.78}^{+1.81}$ $609.91_{-5.02}^{+5.21}$ $601.81_{-2.58}^{+2.51}$ $731.18_{-12.48}^{+8.95}$
$\chi^{2}$/dof - 14.78/34 27.25/13 15.94/13 75.76/19
Figure 6: The inferred viewing angle $\chi$ of the line of sight relative to the magnetar’s spin axis. The dashed line and shaded band show the mean and standard deviation of the four independent measurements.
@@ III Discussion
The distance measurements of the fireballs to the magnetar surface provide a very stringent constraint on the trigger mechanism of the bursts. Generally speaking, highly twisted magnetic fields in the magnetosphere of magnetars are very likely to extend from the stellar interior, where the toroidal component could dominate. In other words, the energy release of magnetars is ultimately a result of the diffusion of magnetic helicity from the interior to the exterior. The connection between the internal and external fields exerts vast stress on the solid crust of magnetar. Then, along with the spin-down of magnetar, the shear stress is gradually accumulated to finally exceed the yield limit of the crust, leading to plastic deformation of the crust and triggering a starquake. In principle, the released strain energy can heat the deforming crustal region, leading to the formation of a hot spot and even a fireball trapped by local magnetic field loops near the surface.
@@
Furthermore, we can envision the following physical scenario. As a magnetic reconnection occurs in the magnetosphere, an energy of about $E_{\rm iso}\sim 10^{41}$ erg can be released suddenly. Then, a great number of electron and positron pairs can be excited in the reconnection zone and flow into a larger-scale region, forming a magnetically trapped fireball. The number density of the fireball can be given by (Landau and Lifshitz, 1980; Thompson and Duncan, 1995)
@@
where $e$ and $m_{\rm e}$ are respectively the charge and mass of electrons, $c$ is the speed of light. The internal temperature of the fireball $T_{\rm in}$ can be roughly estimated to be $k_{\rm B}T_{\rm in}=k_{\rm B}({3E_{\rm iso}/4\pi l^{3}a})^{1/4}=68(E_{\rm iso}/10^{41}{\rm erg})^{1/4}(l/20{\rm km})^{-3/4}\rm keV$ with $a$ being the radiation constant. The fireball expansion could be quickly suppressed if the pressure of the surrounding large-scale magnetic fields can finally exceed the thermal pressure of the fireball, which requires $B\gtrsim 1.6\times 10^{11}\>({l}/{20\>\mathrm{km}})^{-3/2}({E_{\mathrm{iso}}}/{10^{41}\>\mathrm{erg}})^{1/2}\>\mathrm{G}$ (Thompson and Duncan, 1995). By using a simple dipolar field configuration, the field strength at the distance $d$ can be estimated to $B(d)\sim 3.0\times 10^{12}({d}/{50\>\mathrm{km}})^{-3}\mathrm{G}$, which is obviously strong enough to stop the fireball expansion, where a reference stellar radius $R_{\mathrm{ns}}=12$ km is adopted and the surface field strength is taken as $B_{\mathrm{p}}=2.2\times 10^{14}\>\mathrm{G}$ (Israel et al., 2016). Therefore, the fireball can finally be trapped by the magnetic fields and suspended in the magnetosphere for a time interval.
@@
The rationality of a suspended fireball may also be understood from the perspective of interpreting the key properties of the CRSF absorption, including its central energy, line width, and absorption depth: (i) With the magnetic field strength at a reference distance of 50 km, we can determine the position of the first harmonic of CRSFs in the spectra of the bursts as: $E_{\mathrm{cyc}}=\hbar{eB/m_{\rm e}c}=35.3\left({d/50\ {\rm km}}\right)^{-3}\rm keV$, which is valid for the case of $B\ll B_{\rm c}=4.4\times 10^{13}$ G with $B_{\rm c}$ being the critical value for Landau quantization. Although very simple, this estimate of $E_{\mathrm{cyc}}$ can be very consistent with the values obtained from the spectral fittings, further indicating the self-consistency of the model. However, strictly speaking, some inconsistencies still exist, since the fireball distances inferred from the light curve fittings (see Table 1) are actually somewhat different from the reference value. Here, the crux of the problem may lie in the complication of the magnetic field configuration of the magnetar, which distinguishes itself from the traditional dipolar scenario and complicates the determination of both $B(d)$ and $E_{\rm cyc}$ at a given distance. (ii) It is expected that the widths of CRSF are at least larger than that determined by the Doppler effect as (Meszaros and Nagel, 1985; Harding and Daugherty, 1991; Nishimura, 2003; Schönherr et al., 2007): $\sigma=({k_{\rm B}T/m_{\rm e}c^{2}})^{1/2}E_{\rm cyc}|\cos\theta|=4.9{\rm keV}\left({k_{\rm B}T/{10\rm keV}}\right)^{1/2}\left({E_{\rm cyc}/35\rm keV}\right)|\cos\theta|$, where $\theta$ is the angle between the photon direction and the magnetic field vector. This value is smaller than, but of the same order of magnitude as, the CRSF widths derived from the spectral fittings. However, since the fireball is actually distributed within a spatial range of a few tens of kilometers, the magnetic field strength in the vicinity of the fireball should at least vary on an amplitude of $(d+l)/(d-l)]^{3}$. Given this significant variation in the field strength, it would be not surprising to observe a CRSF line a few times broader than the Doppler broadening. (iii) In the outer region adjacent to the fireball surface where the CRSF could be formed, the column density of electrons and positrons can be estimated by $\mu n_{\rm GJ}\Delta$, where $\mu$ is the multiplicity, $n_{\rm GJ}=\Omega B/(2\pi ce)$ is the Goldreich-Julian density, and $\Delta$ is the thickness of this line-formation region. Then, following Wang et al. (1993), the optical depth of the CRSF absorption can be expressed as $\tau_{\mathrm{CRSF}}=0.3\left({\mu/10^{3}}\right)\left({\Delta/5~\rm km}\right)\left({k_{\rm B}T/\rm 10~keV}\right)^{-1/2}$, which could provide a basic explanation for the optical depths derived from the spectral fittings.
@@
where $\lambda=1/(\sigma_{\rm T}n_{\pm})$ is the mean free path of the thermal photons with $\sigma_{\rm T}$ being the Thomson cross section and $l=20$ km is adopted for giving the numerical coefficient. However, this timescale is very sensitive to the determination of the internal temperature of the fireball because of the presence of the exponential term. Specifically, by solving equation $t_{\rm diff}\sim 1$ s, we can obtain $k_{\rm B}T_{\rm in}=32.7$ keV for $l=20$ km, which is lower than that inferred from the average internal energy density. Here, it should be noted that the actual temperature within the fireball must exhibit a decreasing distribution from the inside out, which is determined by thermal transfer within the fireball. Given this, it is not surprising to obtain different temperatures from different estimation methods, not to mention that these approaches are rather crude. Then, undoubtedly, a precise calculation of the burst duration requires a detailed characterization of the thermal diffusion.
In addition, more complexity can arise from the unclear trigger mechanism of the bursts. Given the complication of locally twisted magnetic fields, magnetic reconnection is likely to occur repeatedly and with complicated dynamical behavior rather than impulsively as supposed above. In this case, the pressure balance of the fireball cannot keep steady and, thus, an evolving and even multi-peaked light curve rather than a plateau would be obtained. Such complicated bursts can be easily found from the longest bursts of SGR J1935+2154, with a larger number than plateau bursts. Since we cannot know the dynamical evolution of the fireballs of these complicated bursts, it is actually impossible to identify eclipse candidates from them, even though their spectral properties are not obviously different from the plateau ones.
@@ IV Summary
In this paper, we identified one representative plateau-like burst among the long bursts of SGR J1935+2154 detected by GECAM and Fermi/GBM, and found four candidate bursts showing eclipse-like features. We established the co-rotating fireball geometric eclipse model and fitted the light curves of four eclipse bursts. We derived the viewing angle related to the spin axis with $\chi\approx 17^{\circ}$, and the fireball is located at several NS radii away from the magnetar. This distance indicates that the fireball is suspended in the magnetosphere, where the magnetic field strength is estimated to be $B\sim 3\times 10^{12}~\mathrm{G}$. Moreover, we find CRSF features in the spectra of X-ray bursts, which also implies a magnetic field strength similar to that inferred from the light-curve distances. These results not only provide smoking-gun evidence of the co-rotating fireball, which accounts for the energetic X-ray emission during the magnetar activity, but also support an origin of the fireball from magnetic reconnection in the magnetosphere rather than directly from a crustal starquake. We also note that this new eclipse-based method provides a novel tool to constrain the viewing angle and could be applied to other magnetar bursts.
@@ Appendix A Data Reduction
Gravitational wave high-energy Electromagnetic Counterpart All-sky Monitor (GECAM, Li et al., 2021b; Xiao et al., 2022) is originally a constellation of two small X-ray and gamma-ray all-sky observatories funded by the Chinese Academy of Sciences. These two microsatellites, denoted as GECAM-A and GECAM-B, were launched in December 2020. Each GECAM satellite is equipped with 25 Gamma-ray detectors (GRDs) and 8 Charged Particle Detectors (CPDs). Each GRD operates in two readout channels: high gain (HG, 20–300 keV) and low gain (LG, 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The time resolution of GRDs is 0.1 $\mu\rm s$. We select detectors with detector-to-source angles of $\leqslant 50^{\circ}$ in our analysis. Given the dominated emission energy range of magnetar bursts ($\lesssim$100 keV), we focus on using the HG channel for data analysis. The GECAM data reduction tool is GECAMTools (zhang, 2024). Considering the lanthanum absorption edge at 38.89 keV for $\mathrm{LaBr_{3}}$, as this feature potentially leads to systematic uncertainties in spectra, we have ignored the data in the range of 35-40 keV (He et al., 2023).
Fermi Gamma-ray Burst Monitor (GBM, Meegan et al., 2009) has a continuous broadband energy coverage which consists of 12 NaI detectors (8-1000 keV) and 2 Bismuth Germanate (BGO) detectors (0.2-40 MeV). GBM provides three kinds of science data: CSPEC with continuous high energy resolution (1024 ms and 128 energy channels), CTIME with continuous high time resolution (64 ms and 8 energy channels), and TTE (time tagged photon event) with temporal resolution 2 $\mu s$ and energy resolution 128 channels. We choose the detectors with good detector-to-source angles ($\leqslant 50^{\circ}$), and hence the highest signal-to-noise ratio Collazzi et al. (2015); Lin et al. (2020b). The data reduction tool used to generate light curves and spectral data is GBM Data Tools (Goldstein et al., 2022). Response files are generated by GBM Response Generator v1.13 for spectral analysis. Similar to GECAM, we excluded 30-40 keV data for NaI detectors to avoid the iodine K-edge at 33.17 keV (Bissaldi et al., 2009).
@@ Appendix B Spectral Analysis
We process the GECAM event-by-event (EVT) data and GBM TTE data to generate spectral data for 50 ms time-resolved spectral analysis using a data reduction tool. Background spectra are generated from data events in pre- and post-burst time intervals (from $t_{0}$-10 s to $t_{0}$-5 s, and from $t_{0}$+5 s to $t_{0}$+10 s, where $t_{0}$ is the burst time), and fitted with a polynomial function to model the background during the burst. Subsequently, this background would be subtracted from the total spectrum in the spectral fitting process. The raw light curves of the five selected bursts, with a 5 ms time resolution and the estimated background level, are shown in Figure 7. These light curves are very unlikely to result from the simple superposition of multiple pulses, although some nontrivial substructures may still exist.
The software elisa is employed for exploring and analyzing the spectral data (Xue et al., 2025). In performing Bayesian fitting, the likelihood based on the $\chi^{2}$-statistic is used to compute the posterior distribution, and spectral channels are grouped to ensure a minimum of one count per channel. Furthermore, to evaluate the performance of different spectral models, we used two commonly adopted information criteria: the Akaike Information Criterion (AIC) and the Bayesian Information Criterion (BIC) (Liddle, 2007). Both criteria balance the goodness of fit and the complexity of the model, penalizing models with more free parameters. Lower AIC or BIC values indicate a more statistically favored model.
First of all, we perform spectral fits for the representative plateau of Burst A, by using four different spectral models. The first model is a single black body (BB), which is written as
@@
However, only with a single BB, there is always a significant dip within the energy range from $\sim$20 to $\sim$50 keV, as shown in Figure 8. In order to reduce these discrepancies between the spectral data and model, we further consider another two spectral models by invoking an extra spectral component, which is either another BB or a cyclotron resonant scattering feature (CRSF) absorption. Here, the CRSF with a Gaussian profile implemented as a multiplicative model is given by:
@@
where $E_{\mathrm{cyc}}$ is the central energy of the line in units of keV, $\sigma$ is the line width, and $\tau_{\rm CRSF}$ is the optical depth at the line center. Following the same approach as above, the BB temperatures are tied for the time intervals in the plateau phase and the CRSF parameters are tied for all intervals. The results presented in Figure 8 show that that both of these scenarios can help to eliminate the dip around $\sim 35$ keV. Furthermore, the AIC and BIC values listed in Table 2 strongly suggest that the best fit can be provided by the BB + CRSF model. The flux data presented in the left panel of Figure 3 for Burst A is just obtained with this model.
Finally, we further apply the BB + CRSF model to fit the time-resolved spectra of all other selected bursts. In fact, we have also examined the other three spectral models, which are however disfavored by their AIC and BIC values as listed in Table 2. So, for simplicity, we do not present these alternative fitting results. Bursts B-E fittings with the BB + CRSF model are shown in Figure 9. These results demonstrate that the peaks of the time-resolved spectra can always be concentrated around a constant value, and the temporal evolution of the spectra is solely determined by the variation of the emission area. Figure 4 presents the derived fluxes and emission areas versus time. The comparison given in the middle panels of this figure further demonstrates that these spectrum-determining emission areas can agree well with the area variation curve derived from the eclipse modeling.

### v3

@@ II.2 Eclipse Modeling
$\chi$ (deg) [0–180] $17.46_{-6.87}^{+8.99}$ $17.11_{-8.94}^{+9.69}$ $18.54_{-8.16}^{+15.65}$ $8.04_{-3.56}^{+5.39}$
$\xi$ (deg) [0–180] $155.92_{-11.71}^{+9.35}$ $162.22_{-9.20}^{+7.31}$ $149.38_{-23.41}^{+13.22}$ $172.00_{-5.43}^{+3.50}$
$\phi$ (deg) [0–360] $343.75_{-1.00}^{+1.09}$ $297.60_{-0.32}^{+0.33}$ $316.85_{-0.69}^{+0.66}$ $274.27_{-1.21}^{+1.15}$
$d/R_{\mathrm{ns}}$ [1–50] $5.38_{-1.58}^{+3.19}$ $7.65_{-2.53}^{+5.01}$ $8.23_{-3.15}^{+5.89}$ $7.96_{-3.09}^{+6.07}$
$F_{\mathrm{p}}$ (photons cm-2 s-1) [1–1000] $727.62_{-1.78}^{+1.81}$ $609.91_{-5.02}^{+5.21}$ $601.81_{-2.58}^{+2.51}$ $731.18_{-12.48}^{+8.95}$
$\chi^{2}$/dof 14.78/34 27.25/13 15.94/13 75.76/19
Figure 6: The inferred viewing angle $\chi$ of the line of sight relative to the magnetar’s spin axis. The dashed line and shaded band show the mean and standard deviation, respectively, of the four independent measurements.
@@ III Discussion
The distance measurements of the fireballs to the magnetar surface provide a very stringent constraint on the trigger mechanism of the bursts. Generally speaking, highly twisted magnetic fields in the magnetosphere of magnetars are very likely to extend from the stellar interior, where the toroidal component could dominate. In other words, the energy release of magnetars is ultimately a result of the diffusion of magnetic helicity from the interior to the exterior. The connection between the internal and external fields exerts vast stress on the solid crust of magnetars. Then, along with the spin-down of magnetars, the shear stress is gradually accumulated to finally exceed the yield limit of the crust, leading to plastic deformation of the crust and triggering a starquake. In principle, the released strain energy can heat the deforming crustal region, leading to the formation of a hot spot and even a fireball trapped by local magnetic field loops near the surface.
@@
Furthermore, we can envision the following physical scenario: As a magnetic reconnection occurs in the magnetosphere, an energy of about $E_{\rm iso}\sim 10^{41}$ erg can be released suddenly. Then, a great number of electron and positron pairs can be excited in the reconnection zone and flow into a larger-scale region, forming a magnetically trapped fireball. The number density of the fireball can be given by (Landau and Lifshitz, 1980; Thompson and Duncan, 1995)
@@
where $e$ and $m_{\mathrm{e}}$ are the charge and mass of electrons, respectively, and $c$ is the speed of light. The internal temperature of the fireball $T_{\rm in}$ can be roughly estimated to be $k_{\rm B}T_{\rm in}=k_{\rm B}({3E_{\rm iso}/4\pi l^{3}a})^{1/4}=68(E_{\rm iso}/10^{41}{\rm erg})^{1/4}(l/20{\rm km})^{-3/4}\rm keV$ with $a$ being the radiation constant. The fireball expansion could be quickly suppressed if the pressure of the surrounding large-scale magnetic fields can finally exceed the thermal pressure of the fireball, which requires $B\gtrsim 1.6\times 10^{11}\>({l}/{20\>\mathrm{km}})^{-3/2}({E_{\mathrm{iso}}}/{10^{41}\>\mathrm{erg}})^{1/2}\>\mathrm{G}$ (Thompson and Duncan, 1995). By using a simple dipolar field configuration, the field strength at the distance $d$ can be estimated as $B(d)\sim 3.0\times 10^{12}({d}/{50\,\mathrm{km}})^{-3}\,\mathrm{G}$, which is obviously strong enough to stop the fireball expansion, where a reference stellar radius $R_{\mathrm{ns}}=12$ km is adopted and the surface field strength is taken as $B_{\mathrm{p}}=2.2\times 10^{14}\>\mathrm{G}$ (Israel et al., 2016). Therefore, the fireball can finally be trapped by the magnetic fields and suspended in the magnetosphere for a time interval.
@@
The rationality of a suspended fireball may also be understood from the perspective of interpreting the key properties of the CRSF absorption, including its central energy, line width, and absorption depth: (i) With the magnetic field strength at a reference distance of 50 km, we can determine the position of the first harmonic of CRSFs in the spectra of the bursts as $E_{\mathrm{cyc}}=\hbar{eB/m_{\rm e}c}=35.3\left({d/50\ {\rm km}}\right)^{-3}\rm keV$, which is valid for the case of $B\ll B_{\mathrm{c}}=4.4\times 10^{13}\,\mathrm{G}$ with $B_{\mathrm{c}}$ being the critical value for Landau quantization. Although very simple, this estimate of $E_{\mathrm{cyc}}$ can be very consistent with the values obtained from the spectral fittings, further indicating the self-consistency of the model. However, strictly speaking, some inconsistencies still exist, since the fireball distances inferred from the light-curve fittings (see Table 1) are actually somewhat different from the reference value. Here the crux of the problem may lie in the complication of the magnetic field configuration of the magnetar, which distinguishes itself from the traditional dipolar scenario and complicates the determination of both $B(d)$ and $E_{\rm cyc}$ at a given distance. (ii) It is expected that the widths of CRSFs are at least larger than that determined by the Doppler effect as (Meszaros and Nagel, 1985; Harding and Daugherty, 1991; Nishimura, 2003; Schönherr et al., 2007) $\sigma=({k_{\rm B}T/m_{\rm e}c^{2}})^{1/2}E_{\rm cyc}|\cos\theta|=4.9{\rm keV}\left({k_{\rm B}T/{10\rm keV}}\right)^{1/2}\left({E_{\rm cyc}/35\rm keV}\right)|\cos\theta|$, where $\theta$ is the angle between the photon direction and the magnetic field vector. This value is smaller than, but of the same order of magnitude as, the CRSF widths derived from the spectral fittings. However, since the fireball is actually distributed within a spatial range of a few tens of kilometers, the magnetic field strength in the vicinity of the fireball should at least vary on an amplitude of $(d+l)/(d-l)]^{3}$. Given this significant variation in the field strength, it would be not surprising to observe a CRSF line a few times broader than the Doppler broadening. (iii) In the outer region adjacent to the fireball surface where the CRSF could be formed, the column density of electrons and positrons can be estimated by $\mu n_{\rm GJ}\Delta$, where $\mu$ is the multiplicity, $n_{\rm GJ}=\Omega B/(2\pi ce)$ is the Goldreich–Julian density, and $\Delta$ is the thickness of this line formation region. Then, following Wang et al. (1993), the optical depth of the CRSF absorption can be expressed as $\tau_{\mathrm{CRSF}}=0.3\left({\mu/10^{3}}\right)\left({\Delta/5~\rm km}\right)\left({k_{\rm B}T/\rm 10~keV}\right)^{-1/2}$, which could provide a basic explanation for the optical depths derived from the spectral fittings.
@@
where $\lambda=1/(\sigma_{\rm T}n_{\pm})$ is the mean free path of the thermal photons, where $\sigma_{\rm T}$ is the Thomson cross section and $l=20\,\mathrm{km}$ is adopted for giving the numerical coefficient. However, this timescale is very sensitive to the determination of the internal temperature of the fireball because of the presence of the exponential term. Specifically, by solving equation $t_{\rm diff}\sim 1$ s, we can obtain $k_{\rm B}T_{\rm in}=32.7$ keV for $l=20\,\mathrm{km}$, which is lower than that inferred from the average internal energy density. Here it should be noted that the actual temperature within the fireball must exhibit a decreasing distribution from the inside out, which is determined by thermal transfer within the fireball. Given this, it is not surprising to obtain different temperatures from different estimation methods, not to mention that these approaches are rather crude. Then, undoubtedly, a precise calculation of the burst duration requires a detailed characterization of the thermal diffusion.
In addition, more complexity can arise from the unclear trigger mechanism of the bursts. Given the complication of locally twisted magnetic fields, magnetic reconnection is likely to occur repeatedly and with complicated dynamical behavior rather than impulsively as supposed above. In this case, the pressure balance of the fireball cannot remain steady, and thus an evolving and even multipeaked light curve rather than a plateau would be obtained. Such complicated bursts can be easily found from the longest bursts of SGR J1935+2154, with a larger number than plateau bursts. Since we cannot know the dynamical evolution of the fireballs of these complicated bursts, it is actually impossible to identify eclipse candidates from them, even though their spectral properties are not obviously different from the plateau ones.
@@ IV Summary
In this paper, we identified one representative plateau-like burst among the long bursts of SGR J1935+2154 detected by GECAM and Fermi/GBM, and found four candidate bursts showing eclipse-like features. We established the corotating fireball geometric eclipse model and fitted the light curves of four eclipse bursts. We derived the viewing angle related to the spin axis with $\chi\approx 17^{\circ}$, and the fireball is located at several neutron star radii away from the magnetar. This distance indicates that the fireball is suspended in the magnetosphere, where the magnetic field strength is estimated to be $B\sim 3\times 10^{12}~\mathrm{G}$. Moreover, we find CRSF features in the spectra of X-ray bursts, which also implies a magnetic field strength similar to that inferred from the light-curve distances. These results not only provide smoking-gun evidence of the corotating fireball, which accounts for the energetic X-ray emission during the magnetar activity, but also support an origin of the fireball from magnetic reconnection in the magnetosphere rather than directly from a crustal starquake. We also note that this new eclipse-based method provides a novel tool to constrain the viewing angle and could be applied to other magnetar bursts.
@@ Appendix A Data Reduction
GECAM (X. Li et al., 2021b; Xiao et al., 2022) is originally a constellation of two small X-ray and gamma-ray all-sky observatories funded by the Chinese Academy of Sciences. These two microsatellites, denoted as GECAM-A and GECAM-B, were launched in 2020 December. Each GECAM satellite is equipped with 25 gamma-ray detectors (GRDs) and 8 charged particle detectors (CPDs). Each GRD operates in two readout channels, high gain (HG; 20–300 keV) and low gain (LG; 300 keV–5 MeV), which are independent in terms of data processing, transmission, and dead time. The time resolution of GRDs is 0.1 $\mu\rm s$. We select detectors with detector-to-source angles of $\leqslant 50^{\circ}$ in our analysis. Given the dominated emission energy range of magnetar bursts ($\lesssim$100 keV), we focus on using the HG channel for data analysis. The GECAM data reduction tool is GECAMTools (zhang, 2024). Considering the lanthanum absorption edge at 38.89 keV for $\mathrm{LaBr_{3}}$, as this feature potentially leads to systematic uncertainties in spectra, we have ignored the data in the range of 35–40 keV (He et al., 2023).
Fermi/GBM (Meegan et al., 2009) has a continuous broadband energy coverage that which consists of 12 NaI detectors (8–1000 keV) and 2 bismuth germanate detectors (0.2–40 MeV). GBM provides three kinds of science data: CSPEC with continuous high energy resolution (1024 ms and 128 energy channels), CTIME with continuous high time resolution (64 ms and 8 energy channels), and time-tagged photon event (TTE) with temporal resolution of $2\,\mu\mathrm{s}$ and energy resolution of 128 channels. We choose the detectors with good detector-to-source angles ($\leqslant 50^{\circ}$), and hence the highest signal-to-noise ratio Collazzi et al. (2015); Lin et al. (2020b). The data reduction tool used to generate light curves and spectral data is GBM Data Tools (Goldstein et al., 2022). Response files are generated by GBM Response Generator v1.13 for spectral analysis. Similar to GECAM, we excluded 30–40 keV data for NaI detectors to avoid the iodine K edge at 33.17 keV (Bissaldi et al., 2009).
@@ Appendix B Spectral Analysis
We process the GECAM event-by-event (EVT) data and GBM TTE data to generate spectral data for 50 ms time-resolved spectral analysis using a data reduction tool. Background spectra are generated from data events in pre- and post-burst time intervals (from $t_{0}-10\,\mathrm{s}$ to $t_{0}-5\,\mathrm{s}$, and from $t_{0}+5\,\mathrm{s}$ to $t_{0}+10\,\mathrm{s}$, where $t_{0}$ is the burst time) and fitted with a polynomial function to model the background during the burst. Subsequently, this background would be subtracted from the total spectrum in the spectral fitting process. The raw light curves of the five selected bursts, with a 5 ms time resolution and the estimated background level, are shown in Figure 7. These light curves are very unlikely to result from the simple superposition of multiple pulses, although some nontrivial substructures may still exist.
The software elisa is employed for exploring and analyzing the spectral data (Xue et al., 2025). In performing Bayesian fitting, the likelihood based on the $\chi^{2}$-statistic is used to compute the posterior distribution, and spectral channels are grouped to ensure a minimum of one count per channel. Furthermore, to evaluate the performance of different spectral models, we used two commonly adopted information criteria: the AIC and the BIC (Liddle, 2007). Both criteria balance the goodness of fit and the complexity of the model, penalizing models with more free parameters. Lower AIC or BIC values indicate a more statistically favored model.
First of all, we perform spectral fits for the representative plateau of Burst A, by using four different spectral models. The first model is a single BB, which is written as
@@
However, only with a single BB is there always a significant dip within the energy range from $\sim$20 to $\sim$50 keV, as shown in Figure 8. In order to reduce these discrepancies between the spectral data and model, we further consider another two spectral models by invoking an extra spectral component, which is either another BB or a CRSF absorption. Here the CRSF with a Gaussian profile implemented as a multiplicative model is given by
@@
where $E_{\mathrm{cyc}}$ is the central energy of the line in units of keV, $\sigma$ is the line width, and $\tau_{\rm CRSF}$ is the optical depth at the line center. Following the same approach as above, the BB temperatures are tied for the time intervals in the plateau phase and the CRSF parameters are tied for all intervals. The results presented in Figure 8 show that that both of these scenarios can help to eliminate the dip around $\sim 35$ keV. Furthermore, the AIC and BIC values listed in Table 2 strongly suggest that the best fit can be provided by the BB + CRSF model. The flux data presented in the left panel of Figure 3 for Burst A are just obtained with this model.
Finally, we further apply the BB + CRSF model to fit the time-resolved spectra of all other selected bursts. In fact, we have also examined the other three spectral models, which are, however, disfavored by their AIC and BIC values as listed in Table 2. So, for simplicity, we do not present these alternative fitting results. Bursts B–E fittings with the BB + CRSF model are shown in Figure 9. These results demonstrate that the peaks of the time-resolved spectra can always be concentrated around a constant value, and the temporal evolution of the spectra is solely determined by the variation of the emission area. Figure 4 presents the derived fluxes and emission areas versus time. The comparison given in the middle panels of this figure further demonstrates that these spectrum-determining emission areas can agree well with the area variation curve derived from the eclipse modeling.